OpenStudy (anonymous):

3x - 5y = -9 8x - y = 13 Choose one answer. a. (-2,3) b. (3,-2) c. (3,2) d. (2,3)

OpenStudy (anonymous):

See, here multiply second equation with 5 both the sides, what you got??

OpenStudy (anonymous):

i dnt understand

OpenStudy (anonymous):

See, in first equation, what is the coefficient of y??

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Is it??

OpenStudy (anonymous):

I think -5...

OpenStudy (anonymous):

ohhhh yes -5

OpenStudy (anonymous):

And in second equation??

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

rite?

OpenStudy (anonymous):

Firstly make both the y coefficients equal.. How can you do that??

OpenStudy (anonymous):

Yes you are right...

OpenStudy (anonymous):

uhmm idk

OpenStudy (anonymous):

See, y coefficient is -1.. What number I should multiply with -1 to get -5???

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Yes, that is why I am multiply second equation with 5.. Getting??

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

So, multiply both sides of second equation by 5 what will you get??

OpenStudy (anonymous):

idk if this is rite but 8x-5y=65?

OpenStudy (anonymous):

You are close.. See, what I am saying is : 5(8x - y = 13) You did not multiply 5 with 8x that is the mistake..

OpenStudy (anonymous):

ohk so im distributing its 40x-5y=65

OpenStudy (anonymous):

Yes now you are right.. So, what are the two equations now become/ I am writing both the equations: 3x - 5y = -9 ------------1 40x - 5y = 65 ------------2 Now simply subtract them and tell what will you get??

OpenStudy (anonymous):

then i get -37x=-74 x=2?

OpenStudy (anonymous):

Yes you did that very right.. Can you find y now??

OpenStudy (anonymous):

y=1?

OpenStudy (anonymous):

Now solve it again..

OpenStudy (anonymous):

ohk

OpenStudy (anonymous):

y=3

OpenStudy (anonymous):

See, 3x - 5y = -9 Or, 3(2) - 5y = -9 6 - 5y = -9 -5y = -9 - 6 5y = 15 And get y from here..

OpenStudy (anonymous):

Yes, you are right then what is the answer then??

OpenStudy (anonymous):

d!

OpenStudy (anonymous):

That is correct...

OpenStudy (anonymous):

thanks again!!